It took a car driver 30 mins to travel a distance of 50km. What is the average speed?
1
Expert's answer
2021-09-09T17:21:05-0400

"speed=\frac{distance}{time}=\frac{50km}{30min}\times\frac{60min}{1hr}=100\frac{km}{hr}"


Answer: 100 km/hr